# SelfDefend: LLMs Can Defend Themselves against Jailbreaking in a Practical Manner

### v2

@@ 5.3 Measurement Results
Baseline Performance. We first measure the baseline performance of using GPT-3.5/4 only under the tested jailbreaks. As shown in Table 3, these jailbreaks are much more effective on GPT-3.5 compared to GPT-4. The ASR for GPT-3.5 can be as high as 0.256 to 0.980 (average: 0.657), whereas the ASR for GPT-4 typically ranges from 0.047 to 0.330, except for indirect jailbreaks where GPT-4 also exhibits a high ASR (0.74 for DrAttack and 0.900 for Puzzler).
Defense Effectiveness. We then compare ASR after equipping the baseline model with SelfDefend and report additional ASR findings when using only the shadow stack for ablation exploration. For GPT-3.5, both SelfDefend versions—supported by $P_{direct}$ and $P_{intent}$—reduce ASRs but show varying degrees of defense enhancement. SelfDefend with $P_{direct}$ exhibits a modest reduction, with its shadow stack performing worse than GPT-3.5 itself; for instance, its ASR on AutoDAN is 96%, compared to GPT-3.5’s 90%. In contrast, SelfDefend with $P_{intent}$ significantly outperforms the $P_{direct}$ version, achieving a reduction factor ranging from 97.26% to 8.97%.
@@
[…]
where $d_{total}$ denotes the total delay to SelfDefend-protected LLMs, and $d_{normal}$ represents the separate time cost of the target LLM in the normal stack. The metric of $\Delta d$ is essential for normal prompts because it would directly affect the experience of normal users. Figure 3(a) plots the cumulative distribution function (CDF) of $\Delta d$ for normal prompts. We can see that over 95% of the cases incur zero delays across three configurations of SelfDefend, demonstrating that SelfDefend is capable of defending against jailbreaks at the cost of negligible delay for normal users. The only exception occurs with SelfDefend using $P_{intent}$ and GPT-3.5, where around 80% of the cases have no extra delay.
For jailbreak prompts, we also calculate their average $\Delta d$ for different attack scenarios in Figure 3(b). We observe that GPT-3.5’s average $\Delta d$ is under 0.1 seconds for all eight scenarios, while GPT-4 incurs negligible $\Delta d$ only in the normal scenario of AlpacaEval. This is likely because the baseline GPT-4 has good defense performance for most jailbreaks, with its original output already short and the delay minimal.
Figure 4: The training procedure for fine-tuning our open-source defense model.
@@ 6.3 LoRA Fine-Tuning
Table 4: Jailbreak ASR for various defense methods. For ICD, SmoothLLM and SafeDecoding, we present the performance of their enhanced models. For detection-based Perplexity Filter and Llama Guard (2), we report ASRs only on their detection modules. Since SafeDecoding works for a white-box target model, we show its results only on the publicly available Llama-2.
DAN GCG AutoDAN PAIR TAP DrAttack Puzzler MultiJail AlpacaEval
GPT-3.5 GPT-3.5 (baseline) 0.256 0.560 0.900 0.720 0.670 0.780 0.980 0.393 0.977
ICD [73] 0.226 0.230 0.840 0.360 0.330 0.750 0.990 0.321 0.960
Perplexity Filter [26] 1.000 0.030 1.000 1.000 1.000 1.000 1.000 1.000 0.994
SmoothLLM [59] 0.238 0.480 0.930 0.650 0.610 0.850 0.990 0.705 0.968
Llama Guard [25] 0.561 0.410 0.580 0.430 0.470 0.970 0.930 0.952 0.996
Llama Guard 2 [66] 0.441 0.140 0.150 0.370 0.360 0.890 0.640 0.559 0.991
$P_{direct}$-tuned Shadow Stack 0.262 0.080 0.070 0.140 0.210 0.780 0.070 0.749 0.968
$P_{direct}$-tuned SelfDefend 0.111 0.060 0.070 0.070 0.170 0.620 0.070 0.302 0.948
$P_{intent}$-tuned Shadow Stack 0.297 0.080 0.090 0.160 0.240 0.180 0.200 0.578 0.996
$P_{intent}$-tuned SelfDefend 0.125 0.050 0.080 0.120 0.200 0.160 0.200 0.260 0.975
Double shadow stack 0.213 0.040 0.050 0.100 0.130 0.180 0.070 0.470 0.966
Double shadow stack+GPT-3.5 0.091 0.030 0.050 0.060 0.100 0.160 0.070 0.187 0.947
GPT-4 GPT-4 (baseline) 0.047 0.080 0.190 0.330 0.310 0.740 0.900 0.076 0.973
ICD [73] 0.062 0.050 0.030 0.230 0.230 0.430 0.640 0.051 0.970
Perplexity Filter [26] 1.000 0.030 1.000 1.000 1.000 1.000 1.000 1.000 0.994
SmoothLLM [59] 0.030 0.070 0.180 0.330 0.330 0.910 0.880 0.651 0.971
Llama Guard [25] 0.561 0.410 0.580 0.430 0.400 0.980 0.930 0.952 0.996
Llama Guard 2 [66] 0.441 0.140 0.150 0.350 0.330 0.910 0.640 0.559 0.991
$P_{direct}$-tuned Shadow Stack 0.262 0.040 0.070 0.170 0.220 0.710 0.120 0.717 0.969
$P_{direct}$-tuned SelfDefend 0.032 0.010 0.050 0.150 0.150 0.580 0.070 0.060 0.947
$P_{intent}$-tuned Shadow Stack 0.284 0.080 0.070 0.190 0.190 0.180 0.190 0.565 0.995
$P_{intent}$-tuned SelfDefend 0.034 0.040 0.060 0.190 0.160 0.170 0.140 0.044 0.970
Double shadow stack 0.198 0.020 0.040 0.120 0.130 0.180 0.120 0.467 0.968
Double shadow stack+GPT-4 0.026 0.010 0.040 0.120 0.110 0.170 0.070 0.038 0.945
Llama-2 Llama-2-7b-chat (baseline) 0.678 0.570 0.680 0.590 0.610 0.880 0.990 0.143 0.988
ICD [73] 0.474 0.700 0.560 0.310 0.320 0.220 1.000 0.946 0.898
SafeDecoding [82] 0.655 0.550 0.740 0.560 0.590 0.640 1.000 0.857 0.981
Perplexity Filter [26] 1.000 0.000 1.000 1.000 1.000 1.000 1.000 1.000 0.994
SmoothLLM [59] 0.681 0.930 0.950 0.820 0.810 0.980 1.000 1.000 0.993
Llama Guard [25] 0.561 0.400 0.580 0.460 0.420 0.840 0.930 0.952 0.996
Llama Guard 2 [66] 0.441 0.170 0.150 0.410 0.350 0.890 0.640 0.559 0.991
(7b-chat) $P_{direct}$-tuned Shadow Stack 0.257 0.060 0.050 0.250 0.190 0.360 0.090 0.737 0.970
$P_{direct}$-tuned SelfDefend 0.214 0.040 0.050 0.220 0.180 0.310 0.090 0.102 0.959
$P_{intent}$-tuned Shadow Stack 0.289 0.110 0.080 0.240 0.140 0.150 0.220 0.587 0.991
$P_{intent}$-tuned SelfDefend 0.242 0.090 0.070 0.210 0.140 0.150 0.220 0.063 0.980
Double shadow stack 0.198 0.050 0.040 0.180 0.140 0.120 0.040 0.479 0.965
Double shadow stack+Llama-2 0.164 0.030 0.040 0.160 0.140 0.110 0.040 0.057 0.954
@@ 7.1 Experimental Setup
Target LLMs, Benchmarks, and Environment. As shown in Table 4, our tuned defense models are paired with protect three popular LLMs: GPT-3.5, GPT-4, and the open-source Llama-2-7b-chat [67]. We use the same benchmark datasets as in Table 2 (see §5). We follow the same procedure and configuration as in §5 to generate the tested jailbreak prompts for GPT-3.5 and GPT-4. The process is similar for testing the target model Llama-2-7b-chat; we set the target model for query-based PAIR, TAP and DrAttack to Llama-2-7b-chat to generate their jailbreak prompts. Our evaluations are implemented using PyTorch 2.2.1 and conducted on an NVIDIA TESLA H800 GPU.
Baselines. We compare our framework with popular jailbreak defense methods, including ICD [73], SafeDecoding[82], Perplexity Filter [26], SmoothLLM [59], Llama Guard [25] and Llama Guard 2 [66]. Specifically, ICD adds in-context demonstrations in input prompts to enhance the safety of the target model. We adopt the same 1-shot demonstration as the ICD in [95]. SafeDecoding seeks to methodically scrutinize safety-related disclaimers and amplify the probabilities of their associated token sequences. We only show the defense performance of SafeDecoding on Llama-2-7b-chat since it requires fine-tuning an expert model based on the target model. Perplexity Filter leverages a Llama-2-7b model to calculate the perplexity of the input prompt. A jailbreak is considered to happen when the perplexity exceeds a threshold. We set this threshold at the maximum perplexity of any prompt in the JailbreakBench dataset of harmful behavior prompts. SmoothLLM perturbs the jailbreak prompts with character-level changes to enable the target LLM to perform defense. In this paper, we set SmoothLLM to conduct character swapping with a 10% perturbation percentage. Llama Guard is a fine-tuned Llama-2-7b model designed to detect the toxicity category of input prompts. Llama Guard 2 is a fine-tuned version of Llama-3-8B [18].
@@ 7.2 Defense Effectiveness
Comparison with Existing Defense Methods. As shown in Table 4, our tuned models under SelfDefend achieve satisfactory defense effects under all types of jailbreaks and deliver state-of-the-art (SOTA) results in most cases, while previous defenses fail to show promising performance under different attacks. For ICD, the tested ASRs under Puzzler are not less than 99% for both GPT-3.5 and Llama-2. For SafeDecoding, it also shows weakness against Puzzler and MultiJail. The Perplexity Filter is only effective against GCG and lacks resistance to other jailbreaks. Although it exhibits slightly better ASR against GCG than our methods for GPT-3.5 and Llama-2, this is because its filter threshold is set to the maximum perplexity among all prompts in JailbreakBench, making it trivial to detect GCG with a garbled suffix. SmoothLLM does not show satisfactory defense performance against DrAttack and Puzzler for all three target models. Its ASRs even exceed 80% on Llama-2 for all attack methods except DAN. Llama Guard exhibits poor jailbreak detection for DrAttack, Puzzler, and MultiJail, while Llama Guard 2 presents unacceptable flaws for DrAttack.
Across all attacks and target models, the average ASR of our $P_{direct}$-tuned shadow stack (27.76%) is 37.61% lower than that of Llama Guard (65.37%) and 16.74% lower than that of Llama Guard 2 (44.50%). Similarly, the average defense performance of our $P_{intent}$-tuned shadow stack (22.46%) outperforms Llama Guard by 42.91% and Llama Guard 2 by 22.04%. Furthermore, SelfDefend achieves ASRs comparable with all other defense methods under AlpacaEval, indicating its compatibility with normal prompts.
@@
$P_{direct}$-tuned SelfDefend v.s. $P_{intent}$-tuned SelfDefend. In general, SelfDefend with the direct prompt performs better than SelfDefend with the intent prompt on human-based, optimization-based, and generation-based jailbreaks, while $P_{intent}$-tuned SelfDefend is more effective against indirect and multilingual attacks. DAN, GCG, AutoDAN, PAIR, and TAP contain explicitly harmful instructions, whereas DrAttack, Puzzler, and MultiJail do not. This phenomenon aligns with our original intent for designing the two prompts, since $P_{direct}$ is used to directly extract harmful parts, and $P_{intent}$ can identify the true intentions hidden by implicit attacks. For example, DrAttack deconstructs the sentence components of the original jailbreak prompt and substitutes them with harmless words, making it challenging for SelfDefend (with $P_{direct}$) to extract any harmful content. Although Puzzler is also an indirect attack, the prompts it generates include offensive content such as inducing clues for jailbreak, resulting in a lower ASR for SelfDefend with $P_{direct}$. Moreover, the false positive rate of the $P_{intent}$-tuned SelfDefend on normal prompts is lower than that of the $P_{direct}$-tuned SelfDefend.
@@
Case Study. Besides the quantitative analysis above, we conduct case studies to understand in depth the advantages and pitfalls of SelfDefend compared to other defenses. Figure 5 illustrates an example of defense methods against a jailbreak prompt. Due to page limitation, readers may refer to Appendix A for its detailed analysis, as well as other cases.
DAN GCG AutoDAN PAIR TAP DrAttack Puzzler MultiJail
$P_{direct}$-tuned model Generated Attack Prompts 0.685 0.851 0.784 0.828 0.827 0.737 0.663 0.677
Identified Harmful Prompts 0.932 0.946 0.939 0.898 0.900 0.785 0.728 0.687
$P_{intent}$-tuned model Generated Attack Prompts 0.686 0.852 0.784 0.831 0.827 0.732 0.667 0.669
Identified Harmful Intentions 0.919 0.909 0.908 0.875 0.874 0.857 0.788 0.766
Figure 6: Average $\Delta d$ introduced by the tuned models on Llama-2-7b-chat.

### v3

@@ 5.3 Measurement Results
Baseline Performance. We first measure the baseline performance of using GPT-3.5/4 only under the tested jailbreaks. As shown in Table 3, these jailbreaks are much more effective on GPT-3.5 compared to GPT-4. The ASR for GPT-3.5 can be as high as 0.256 to 0.980 (average: 0.655), whereas the ASR for GPT-4 typically ranges from 0.047 to 0.330, except for indirect jailbreaks where GPT-4 also exhibits a high ASR (0.74 for DrAttack and 0.900 for Puzzler).
Defense Effectiveness. We then compare ASR after equipping the baseline model with SelfDefend and report additional ASR findings when using only the shadow stack for ablation. For GPT-3.5, both SelfDefend versions—supported by $P_{direct}$ and $P_{intent}$—reduce ASRs but show varying degrees of defense enhancement. SelfDefend with $P_{direct}$ exhibits a modest reduction, with its shadow stack performing worse than GPT-3.5 itself; for instance, its ASR on AutoDAN is 96%, compared to GPT-3.5’s 90%. In contrast, SelfDefend with $P_{intent}$ significantly outperforms the $P_{direct}$ version, achieving a reduction factor ranging from 97.26% to 8.97%.
@@
Figure 3: The CDF plot of $\Delta d$ for normal prompts.
Figure 4: The training procedure for fine-tuning our open-source defense models.
@@
$\Delta d=d_{total}-d_{normal},$ (1)
where $d_{total}$ denotes the total delay to SelfDefend-protected LLMs, and $d_{normal}$ represents the separate time cost of the target LLM in the normal stack. The metric of $\Delta d$ is essential for normal prompts because it would directly affect the experience of normal users. Figure 3 plots the cumulative distribution function (CDF) of $\Delta d$ for normal prompts. We can see that over 95% of the cases incur zero delays across three configurations of SelfDefend, demonstrating that SelfDefend is capable of defending against jailbreaks at the cost of negligible delay for normal users. The only exception occurs with SelfDefend using $P_{intent}$ and GPT-3.5, where around 80% of the cases have no extra delay.
For jailbreak prompts, we also calculate their average $\Delta d$ for different attack scenarios in Figure 5 (see Appendix B). GPT-3.5’s average $\Delta d$ is under 0.1 seconds for all 10 scenarios, while GPT-4 incurs negligible $\Delta d$ only in the normal scenario of AlpacaEval. This is likely because the baseline GPT-4 has good defense performance for most jailbreaks, with its original output already short and the delay minimal.
@@ 7.1 Experimental Setup
Target LLMs, Benchmarks, and Environment. As shown in Table 4, our tuned defense models are tested to protect six popular LLMs: the proprietary GPT-3.5, GPT-4, Claude-3.5-sonnet [6], the open-source Llama-2-chat [69] with 7B and 13B sizes, and Mistral-7B-Instruct-v0.2 [68], covering diverse model architectures and sizes. Due to page limitation, the results for Claude-3.5 and Llama-2-13b-chat are reported in Appendix A. We use the same benchmark datasets as in Table 2 and follow the similar procedures and configurations as in §5 to generate the tested jailbreak prompts. For query-based attacks (RLbreaker, PAIR, TAP, LLM-Fuzzer, and DrAttack), we also generate jailbreaks for the corresponding target LLMs. Our evaluations are implemented using PyTorch 2.2.1 and conducted on an NVIDIA TESLA H800 GPU.
Baselines. We compare our framework with popular jailbreak defense methods, including ICD [75], SafeDecoding[84], Perplexity Filter [27], SmoothLLM [60], Llama Guard [26], Llama Guard 2/3 [67, 19]. Specifically, ICD adds in-context demonstrations in input prompts to enhance the safety of the target model. We adopt the same 1-shot demonstration as the ICD in [98]. SafeDecoding seeks to methodically scrutinize safety-related disclaimers and amplify the probabilities of their associated token sequences. We only show the defense performance of SafeDecoding on Llama-2-7b-chat since it requires fine-tuning an expert model based on the target model. Perplexity Filter leverages a Llama-2-7b model to calculate the perplexity of the input prompt. A jailbreak is considered to happen when the perplexity exceeds a threshold. We set this threshold at the maximum perplexity of any prompt in the JailbreakBench dataset of harmful behavior prompts. SmoothLLM perturbs the jailbreak prompts with character-level changes to enable the target LLM to perform defense. In this paper, we set SmoothLLM to conduct character swapping with a 10% perturbation percentage. Llama Guard is a fine-tuned Llama-2-7b model designed to detect the toxicity category of input prompts. Llama Guard 2 and 3 are similar, but fine-tuned on Llama-3 and Llama-3.1 (8B), respectively.
@@
Table 4: Jailbreak ASR for various defense methods. For ICD and SafeDecoding, we present the performance of their enhanced models. For detection-based Perplexity Filter, SmoothLLM and Llama Guards, we report ASRs only on their detection modules. Since SafeDecoding works for a white-box target model, we show its results on the publicly available Llama-2 and Mistral.
Model DAN GCG AutoDAN RLbreaker PAIR TAP LLM-Fuzzer DrAttack Puzzler MultiJail AlpacaEval
GPT-3.5 GPT-3.5 (baseline) 0.256 0.560 0.900 0.650 0.720 0.670 0.640 0.780 0.980 0.393 0.977
ICD [75] 0.226 0.230 0.840 0.140 0.360 0.330 0.390 0.750 0.990 0.321 0.960
Perplexity Filter [27] 1.000 0.030 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 0.994
SmoothLLM [60] 0.238 0.480 0.930 0.320 0.650 0.610 0.490 0.850 0.990 0.267 0.968
Llama Guard [26] 0.561 0.410 0.580 0.790 0.430 0.470 0.710 0.970 0.930 0.952 0.996
Llama Guard 2 [67] 0.441 0.140 0.150 0.410 0.370 0.360 0.180 0.890 0.640 0.559 0.991
Llama Guard 3 [19] 0.343 0.080 0.130 0.110 0.230 0.290 0.080 0.610 0.420 0.378 0.986
$P_{direct}$-tuned Shadow Stack 0.262 0.080 0.070 0.040 0.140 0.210 0.050 0.780 0.070 0.749 0.968
$P_{direct}$-tuned SelfDefend 0.111 0.060 0.070 0.040 0.070 0.170 0.030 0.620 0.070 0.302 0.948
$P_{intent}$-tuned Shadow Stack 0.297 0.080 0.090 0.050 0.160 0.240 0.040 0.180 0.200 0.578 0.996
$P_{intent}$-tuned SelfDefend 0.125 0.050 0.080 0.050 0.120 0.200 0.030 0.160 0.200 0.260 0.975
Double shadow stack 0.213 0.040 0.050 0.010 0.100 0.130 0.020 0.180 0.070 0.470 0.966
Double shadow stack+GPT-3.5 0.091 0.030 0.050 0.010 0.060 0.100 0.010 0.160 0.070 0.187 0.947
GPT-4 GPT-4 (baseline) 0.047 0.080 0.190 0.290 0.330 0.310 0.190 0.740 0.900 0.076 0.973
ICD [75] 0.062 0.050 0.030 0.010 0.230 0.230 0.050 0.430 0.640 0.051 0.970
Perplexity Filter [27] 1.000 0.030 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 0.994
SmoothLLM [60] 0.030 0.070 0.180 0.040 0.330 0.330 0.220 0.910 0.880 0.048 0.971
Llama Guard [26] 0.561 0.410 0.580 0.660 0.430 0.400 0.700 0.980 0.930 0.952 0.996
Llama Guard 2 [67] 0.441 0.140 0.150 0.340 0.350 0.330 0.150 0.910 0.640 0.559 0.991
Llama Guard 3 [19] 0.343 0.080 0.130 0.090 0.330 0.220 0.110 0.590 0.420 0.378 0.986
$P_{direct}$-tuned Shadow Stack 0.262 0.040 0.070 0.030 0.170 0.220 0.030 0.710 0.120 0.717 0.969
$P_{direct}$-tuned SelfDefend 0.032 0.010 0.050 0.010 0.150 0.150 0.020 0.580 0.070 0.060 0.947
$P_{intent}$-tuned Shadow Stack 0.284 0.080 0.070 0.060 0.190 0.190 0.040 0.180 0.190 0.565 0.995
$P_{intent}$-tuned SelfDefend 0.034 0.040 0.060 0.020 0.190 0.160 0.030 0.170 0.140 0.044 0.970
Double shadow stack 0.198 0.020 0.040 0.010 0.120 0.130 0.010 0.180 0.120 0.467 0.968
Double shadow stack+GPT-4 0.026 0.010 0.040 0.000 0.120 0.110 0.000 0.170 0.070 0.038 0.945
Llama-2 Llama-2-7b-chat (baseline) 0.678 0.570 0.680 0.490 0.590 0.610 0.120 0.880 0.990 0.143 0.988
ICD [75] 0.474 0.700 0.560 0.310 0.310 0.320 0.630 0.220 1.000 0.146 0.898
SafeDecoding [84] 0.655 0.550 0.740 0.630 0.560 0.590 0.610 0.640 1.000 0.857 0.981
Perplexity Filter [27] 1.000 0.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 0.994
SmoothLLM [60] 0.681 0.930 0.950 0.780 0.820 0.810 0.870 0.980 1.000 0.130 0.993
Llama Guard [26] 0.561 0.400 0.580 0.480 0.460 0.420 0.640 0.840 0.930 0.952 0.996
Llama Guard 2 [67] 0.441 0.170 0.150 0.300 0.410 0.350 0.280 0.890 0.640 0.559 0.991
Llama Guard 3 [19] 0.343 0.090 0.130 0.090 0.310 0.280 0.130 0.410 0.420 0.378 0.986
(7b-chat) $P_{direct}$-tuned Shadow Stack 0.257 0.060 0.050 0.020 0.250 0.190 0.020 0.360 0.090 0.737 0.970
$P_{direct}$-tuned SelfDefend 0.214 0.040 0.050 0.020 0.220 0.180 0.010 0.310 0.090 0.102 0.959
$P_{intent}$-tuned Shadow Stack 0.289 0.110 0.080 0.030 0.240 0.140 0.040 0.150 0.220 0.587 0.991
$P_{intent}$-tuned SelfDefend 0.242 0.090 0.070 0.020 0.210 0.140 0.010 0.150 0.220 0.063 0.980
Double shadow stack 0.198 0.050 0.040 0.010 0.180 0.140 0.000 0.120 0.040 0.479 0.965
Double shadow stack+Llama-2 0.164 0.030 0.040 0.010 0.160 0.140 0.000 0.110 0.040 0.057 0.954
Mistral Mistral-7B-Instruct-v0.2 (baseline) 0.685 0.930 0.990 0.410 0.780 0.730 0.450 0.760 0.990 0.276 0.970
ICD [75] 0.679 0.680 0.980 0.430 0.740 0.750 0.810 0.630 0.990 0.286 0.932
SafeDecoding [84] 0.818 0.930 0.970 0.690 0.830 0.780 0.900 0.690 0.990 0.883 0.979
Perplexity Filter [27] 1.000 0.110 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 0.994
SmoothLLM [60] 0.729 0.980 1.000 0.690 0.920 0.850 0.800 0.930 0.990 0.276 0.994
Llama Guard [26] 0.552 0.390 0.990 0.810 0.440 0.410 0.420 0.870 0.930 0.952 0.996
Llama Guard 2 [67] 0.441 0.180 0.110 0.310 0.340 0.350 0.310 0.880 0.620 0.559 0.990
Llama Guard 3 [19] 0.343 0.150 0.140 0.070 0.150 0.290 0.130 0.490 0.420 0.378 0.986
(7B-Instruct $P_{direct}$-tuned Shadow Stack 0.260 0.060 0.050 0.020 0.120 0.220 0.050 0.350 0.120 0.708 0.968
-v0.2) $P_{direct}$-tuned SelfDefend 0.192 0.060 0.050 0.000 0.120 0.210 0.010 0.300 0.120 0.178 0.939
$P_{intent}$-tuned Shadow Stack 0.297 0.070 0.060 0.030 0.100 0.210 0.040 0.060 0.240 0.600 0.993
$P_{intent}$-tuned SelfDefend 0.226 0.070 0.060 0.020 0.080 0.210 0.000 0.060 0.230 0.140 0.964
Double shadow stack 0.208 0.030 0.050 0.010 0.060 0.180 0.030 0.050 0.070 0.498 0.965
Double shadow stack+Mistral 0.151 0.030 0.050 0.000 0.060 0.180 0.000 0.050 0.070 0.121 0.937
@@ 7.2 Defense Effectiveness
Comparison with Existing Defense Methods. As shown in Table 4, our tuned models under SelfDefend achieve satisfactory defense effects under all types of jailbreaks and deliver state-of-the-art (SOTA) results in most cases, while previous defenses fail to show promising performance under different attacks. For ICD, the tested ASRs under Puzzler are not less than 99% for GPT-3.5, Llama-2 and Mistral. For SafeDecoding, it also shows weakness against Puzzler and MultiJail. The Perplexity Filter is only effective against GCG and lacks resistance to other jailbreaks. Although it exhibits slightly better ASR against GCG than our methods for GPT-3.5 and Llama-2, this is because its filter threshold is set to the maximum perplexity among all prompts in JailbreakBench, making it trivial to detect GCG with a garbled suffix. SmoothLLM does not show satisfactory defense performance against DrAttack and Puzzler for all four target models. Llama Guard exhibits poor jailbreak detection for DrAttack, Puzzler, and MultiJail, while Llama Guard 2 presents unacceptable flaws for DrAttack. Llama Guard 3, benefiting from fine-tuning on Llama-3.1-8B, shows better performance than Llama Guard 1 and 2 but is still inferior overall to our fine-tuned defense models.
Across all attacks and target models, the average ASR of our $P_{direct}$-tuned shadow stack (22.03%) is 44.05% lower than that of Llama Guard (66.08%), 19.07% lower than that of Llama Guard 2 (41.10%), and 4.23% lower than that of Llama Guard 3 (26.26%). Similarly, the average defense performance of our $P_{intent}$-tuned shadow stack (18.39%) outperforms Llama Guard by 47.69%, Llama Guard 2 by 22.71%, and Llama Guard 3 by 7.87%. Based on the same base model, Llama-2-7b, our method shows a remarkable improvement over Llama Guard 1. Even though Llama Guard 2 and 3 are fine-tuned on the more advanced Llama-3-8B and Llama-3.1-8B, respectively, our fine-tuned model still outperforms them. This advantage mainly stems from the fact that our approach is more fundamental, extracting harmful objectives from jailbreak prompts, rather than directly assessing the safety of the entire content like the Llama Guard series.
DAN GCG AutoDAN RLbreaker PAIR TAP LLM-Fuzzer DrAttack Puzzler MultiJail
$P_{direct}$-tuned model Generated Attack Prompts 0.685 0.851 0.784 0.732 0.828 0.827 0.673 0.737 0.663 0.677
Identified Harmful Prompts 0.932 0.946 0.939 0.946 0.898 0.900 0.979 0.785 0.728 0.687
$P_{intent}$-tuned model Generated Attack Prompts 0.686 0.852 0.784 0.737 0.831 0.827 0.671 0.732 0.667 0.669
Identified Harmful Intentions 0.919 0.909 0.908 0.899 0.875 0.874 0.901 0.857 0.788 0.766
@@
$P_{direct}$-tuned SelfDefend v.s. $P_{intent}$-tuned SelfDefend. In general, SelfDefend with the direct prompt performs better than SelfDefend with the intent prompt on human-based, optimization-based, and generation-based jailbreaks, while $P_{intent}$-tuned SelfDefend is more effective against indirect and multilingual attacks. This is because DAN, GCG, AutoDAN, RLbreaker, PAIR, TAP, and LLM-Fuzzer contain explicit harmful instructions, whereas DrAttack, Puzzler, and MultiJail do not. This phenomenon aligns with our original intent for designing the two prompts, since $P_{direct}$ is used to directly extract harmful parts, and $P_{intent}$ can identify the true intentions hidden by implicit attacks. For example, DrAttack deconstructs the sentence components of the original jailbreak prompt and substitutes them with harmless words, making it challenging for SelfDefend (with $P_{direct}$) to extract any harmful content. Although Puzzler is also an indirect attack, the prompts it generates include offensive content such as inducing clues for jailbreak, resulting in a lower ASR for SelfDefend with $P_{direct}$. Moreover, the false positive rate of the $P_{intent}$-tuned SelfDefend on normal prompts is lower than that of the $P_{direct}$-tuned SelfDefend.
Certain Outlier Cases. We observe that some defense methods exhibit better performance than our frameworks in certain outlier cases. These cases are typically due to the specific design of the defense methods. For example, Perplexity Filter achieves the lowest ASR against GCG on GPT-3.5 and Llama-2, as it sets the filter threshold to the maximum perplexity among all prompts in JailbreakBench, making it trivial to detect GCG with a garbled suffix. SmoothLLM performs the lowest ASR (0.030) against DAN on GPT-4. Since DAN itself can only achieve an ASR of 0.047 on GPT-4, it is not surprising that SmoothLLM, which votes by accessing the target LLM multiple times, has a slightly lower ASR. Additionally, ICD achieves the lowest ASR against AutoDAN on GPT-4, but we also observe that ICD is much more effective in weakening attacks on GPT-4 than on other target models, which indicates the success of ICD on the powerful GPT-4 by adding in-context demonstrations to enhance the safety of the target model. In addition, AutoDAN adds the harmful goal at the end of the jailbreak prompt, and therefore, ICD shows a strong rejection rate for AutoDAN.
@@
Case Studies. Besides the quantitative analysis above, we conduct case studies to understand in depth the advantages and pitfalls of SelfDefend compared to other defenses. Due to page limitation, readers may refer to Appendix C.